Circular No. 8269

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004C IN NGC 3683 C. C. Dudley and J. Fischer, Naval Research Laboratory, report the

discovery of a new point source (mag K′ ∼ 14) in NGC 3683 observed with the U.S. Naval Observatory 1.55-m telescope (+ 2.1-μm K′ filter + ASTROCAM 1024×1024 ALADDIN InSb imager) at Flagstaff on Jan. 12.5 UT. The position of the source is α = 11^{h}27^{m}29^{s}.72, δ = +56^{o}52′48′′.2 (equinox 2000.0), based on positions of three 2MASS stars, which is 17′′ west and 10′′ north of the K′ nucleus of NGC 3683. The source shows no perceptible drift over 8 min (seeing ∼ 1′′) and was not present on 2003 May 18, the most recent prior observation with ASTROCAM.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004C, obtained by P. Berlind on Jan. 15.51 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ic supernova, similar to SN 1990U ∼ three weeks past maximum (see Matheson et al. 2001, A.J. 121, 1648). The spectrum is highly reddened and shows a strong, narrow Na I D absorption at the velocity of the host galaxy with an equivalent width of 0.3 nm. The dust maps of Schlegel et al. (1998, Ap.J. 500, 525) indicate a low value for Galactic reddening of E(B − V) = 0.015.

COMETS C/2002 X16, 2002 Y4, 2003 A3, 2003 A4, 2003 A5 (SOHO)

Further to IAUC 8266, K. Battams reports measures of additional Kreutz comets found by X. Leprette and M. Oates (C/2002 X16), R. Kracht (C/2002 Y4), and R. Matson (remaining objects) on SOHO website images; C/2002 X16, C/2002 Y4, and C/2003 A4 were visible in C2-coronagraph images, while the 2003 comets were each visible in C3 images. Reduced positions and orbital elements by B. G. Marsden appear on MPEC 2004-A52.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 X16 Dec. 10.621  17^{h}13.^{m}1  −24^{o}48′
C/2002 Y4  26.171  18 25.6  −24 50
Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 A3 Jan. 6.721  19^{h}21.^{m}9  −24^{o}43′
C/2003 A4  14.904  20 10.9  −25 01
C/2003 A5  15.096  20 03.6  −23 54
2004 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green